if the width is 15 shorter than the length and the perimeter is 162 what is the dimension
Answer by Cromlix(4381)

Length = x
Width = x - 15
Perimeter = 2*length + 2*width
162 = 2(x) + 2(x - 15)
162 = 2x + 2x - 30
162 + 30 = 4x
4x = 192
x = 48
Length = 48
Width = 33
Hope this helps.
:-)